OpenStudy (anonymous):

help plzzzz Skylar's grades on four math tests are 85, 78, 77, and 69. What does Skylar need to score on the next test in order to have a mean score of 80? 83 90 91

OpenStudy (anonymous):

\[\frac{85+78+77+69+x}{5}=80\]solve for \(x\)

OpenStudy (anonymous):

\(\Large \color{blue}{\rightarrow x + 85 + 78 + 77 + 69 = ? }\)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Ok look first add up the numbers

OpenStudy (anonymous):

then tell me what you got

OpenStudy (anonymous):

309

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

help me plzz